# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ 1 Introduction
Evaluating large language models against expensive oracle labels (human preferences, expert audits, downstream KPIs) is prohibitively costly at scale. The standard workaround is to use cheap LLM judges as proxies: collect judge scores on every sample, reserve oracle labels for a small validation slice, and hope the correlation is strong enough. This approach is ubiquitous in practice but statistically unsound.
@@
    Preference inversion. Judge scores $S$ are treated as if they were rewards $Y$ on the same scale. They are not. Without calibration, higher $S$ can predict lower $Y$, inverting rankings entirely. The “You’re absolutely right!” phenomenon illustrates the cost: sycophantic responses score high on cheap proxies but low on actual user value. Teams know they should evaluate against long-term outcomes, but oracle labels are expensive, so they settle for proxies that may point the wrong direction.
@@
    Invalid confidence intervals. Naive CIs on uncalibrated judge scores achieve near-0% coverage: their 95% CIs capture the true value in 0/50 seeds. Bootstrap inference with bias-corrected estimation dramatically improves coverage to $\sim$}95% for Direct mode. (Figure 6: at 5% oracle fraction, calibration uncertainty contributes over half of total variance.)
@@
Recent work recognizes that LLM judges require calibration (Lee et al., 2025), but existing methods address only the simplest setting: single-model accuracy with binary outcomes. Guerdan et al. (2025) use doubly robust methods to address external validity when LLM judges are prompted with demographic personas, a problem complementary to our focus on judge calibration for policy ranking. The harder problem of ranking multiple policies with continuous rewards under distribution shift remains inadequately addressed: no existing method provides calibration, overlap diagnostics, and valid uncertainty quantification in a unified framework. Without such a solution, practitioners either pay for full oracle labeling or accept invalid comparisons. CJE fills this gap.
@@
We introduce CJE, a framework that addresses all three failures. CJE casts offline evaluation as calibrated surrogate estimation with three core components:
  • AutoCal-R (reward calibration): Mean-preserving isotonic regression from judge score $S$ to oracle labels $Y$, with an automatic two-stage fallback for covariate-dependent bias (e.g., response length).
  • SIMCal-W (Surrogate-Indexed Monotone Calibration for Importance Weights): Variance-optimal stacking of $S$-monotone, unit-mean weight candidates, stabilizing weights for improved ESS.
  • OUA inference: Bootstrap with augmented inverse-propensity-weighted (AIPW) bias correction ($\hat{\theta}_{\mathrm{aug}}$) that propagates calibration uncertainty into confidence intervals, achieving near-nominal coverage.
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019), learning $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, and applies it to LLM evaluation with auditable transport diagnostics. These components instantiate a single design rule, Design-by-Projection (DbP), that encodes justified knowledge as model restrictions. Building on established semiparametric theory, we show that when justified knowledge defines a restricted statistical model, the efficiency bound in the restricted model is at most the bound in the baseline model (Theorem 2); with cross-fitting, our estimators attain the surrogate information bound.
Beyond variance control, CJE includes a mean transport test: for each target policy $\pi^{\prime}$ or time period, we test whether the mean residual $\mathbb{E}[Y-f(S,X)]$ is zero. This is a necessary condition for reusing a single calibration function across policies or over time; if it fails, surrogate-only evaluation is systematically biased. In our Arena benchmark, the base-trained calibration transports cleanly to clone, premium, and prompt-engineered policies, but fails for an adversarial “unhelpful” policy, where mean residuals reveal a 0.31-0.31 level shift. This test operationalizes a key epistemological shift: rather than assuming transportability holds globally (an untestable identification condition), we treat it as a falsifiable hypothesis testable for each policy. When the test fails, the failure is visible and actionable: either recalibrate with policy-specific data or refuse absolute value claims for that cell. This transforms a brittle identification assumption into a standardized audit.
@@
  • OPE fails unexpectedly. We expected weight stabilization to enable OPE; it did not. Despite SIMCal-W boosting ESS from $1\%$ to $80\%$, calibrated IPS remains near-random (47% pairwise). CLE explains why: high ESS is necessary but not sufficient when the logger rarely visits target-typical regions.
@@
    Methods: AutoCal-R (mean-preserving calibration), SIMCal-W (weight stabilization for improved ESS), and OUA inference (calibration-aware CIs).
@@
    Theory: CJE adapts the surrogate index framework to LLM evaluation with a key innovation: transportability becomes auditable rather than assumed via the mean transport test. Inference is grounded in the efficient influence function for missing-outcome estimation, explaining why bootstrap with $\hat{\theta}_{\mathrm{aug}}$ achieves $\sim$}95% coverage where naive methods achieve 0%. The CLE diagnostic provides a novel explanation for OPE failure despite high ESS, formalizing a precision floor under limited overlap.
@@
Readers primarily interested in deployment can focus on Section 3 (pipeline overview), Section 5.3 (six practical rules), and Table 6 (estimator selection guide).
@@ 2 Background and Setup
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019): learn a mapping $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, then use it for downstream estimation. For policy value estimation, we require only mean transport: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$, which we treat as auditable via a small oracle slice in each target context (3). We additionally impose monotonicity and mean-preservation as Design-by-Projection restrictions. These are not identification assumptions, but deliberate regularization choices that stabilize calibration and prevent preference inversions.
@@
where $\sigma_{\mathcal{T}}^{2}:=\mathrm{ess\,inf}_{(x,a)\in\mathcal{T}}\mathrm{Var}(Y\mid X{=}x,A{=}a)$ is the minimal outcome noise in $\mathcal{T}$, and $\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})$ measures shape mismatch inside $\mathcal{T}$ (proof in Appendix C.12). The diagnostic has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions; (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, inflating the floor even with good coverage; and (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
@@
Automatic judges (LLM-as-judge or preference models) provide scalable scoring (Ouyang et al., 2022; Bai et al., 2022; Zheng et al., 2023; Kim et al., 2024; Kocmi & Federmann, 2023) but are correlational and may drift (Dietz et al., 2025). Viewing $S$ as a surrogate connects to surrogate endpoints and mediation (Prentice, 1989; Robins & Greenland, 1992; Frangakis & Rubin, 2002; Pearl, 2012; VanderWeele, 2015). The surrogate paradox, wherein a policy improves $S$ and $S$ correlates with $Y$ yet the policy harms $Y$, can arise even with high correlation (VanderWeele, 2013); our transport test (3) provides a necessary check. Under mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=\mu(S)$), calibrating $R=f(S)$ preserves $V(\pi^{\prime})=\mathbb{E}[f(S^{\pi^{\prime}})]$ and supplies a one-dimensional index that stabilizes weights.
@@
Figure 1: Judge-as-surrogate under policy and environment shift. $X$: prompt; AA: response; $S$: judge score; $Y$: oracle label; $\pi$: policy; $E$: environment. $S$ and $Y$ are parallel measurements of the response. Solid arrows: causal structure ($\pi\to A$: policy determines response). Dashed red arrows: mechanism-shift edges indicating the $S$-$Y$ calibration may differ across policies or environments; these are not causal effects of $\pi$ or $E$ on $Y$ holding $(X,A)$ fixed. CJE tests whether calibration transports via $\mathbb{E}[Y-f(S,X)]=0$.
@@
OUA uncertainty & IF stacking.
Treating learned $R=\hat{f}(S)$ as fixed understates uncertainty when the oracle slice is small. CJE addresses this via Oracle-Uncertainty-Aware (OUA) inference, which propagates calibration uncertainty into confidence intervals using a delete-one-fold jackknife (Efron & Stein, 1981; Bickel et al., 1993). For ensemble estimation, CJE stacks multiple estimators by minimizing IF covariance over the simplex (Wolpert, 1992; van der Laan et al., 2007). Details in Section 3.
@@ 3 Methods
The CJE pipeline has three components that directly address the three failures identified above: reward calibration to prevent preference inversion, weight stabilization to control IPS variance, and uncertainty-aware inference that dramatically improves CI coverage. Figure 2 provides an overview.
@@
CJE follows one design rule, Design-by-Projection (DbP), applied to each object in the pipeline: (i) calibrate the reward (projection onto a monotone cone), (ii) stabilize ratios (projection onto a unit-mean, $S$-monotone cone), (iii) compute an orthogonalized estimator (projection onto a nuisance-orthogonal subspace), and (iv) optionally hedge variance by stacking (projection onto a simplex). All learners are cross-fitted; by the Efficiency via Model Restriction theorem, these projections encode justified knowledge that lowers the efficiency bound (see Section 4).
3.1 Reward calibration (AutoCal-R: isotonic in $S$ with an automatic two-stage fallback)
@@
  • Monotone mode (default). Isotonic regression on $S$: $\hat{f}_{\uparrow}\in\arg\min_{f\in\mathcal{M}_{\uparrow}}\sum_{i\in O}\big(Y_{i}-f(S_{i})\big)^{2}$. PAVA preserves the slice mean exactly.
  • Two-stage mode (automatic fallback). Fit a smooth index $Z(S,X)=g(S,X)$ (splines+ridge), map to mid-ranks $U=\mathrm{ECDF}\{Z(S,X)\}$, then fit isotonic $\hat{h}_{\uparrow}(U)$. Predictions are $R=\hat{h}_{\uparrow}(\mathrm{ECDF}\{g(S,X)\})$.
Remark (covariate flexibility). The terminal isotonic step preserves mean-honesty regardless of the first-stage model, so $g$ can be any cross-fitted learner (splines, random forests, gradient boosting) and can incorporate covariates $X$ beyond $S$. In our experiments, including response length as a first-stage covariate significantly improves ranking accuracy: LLM judges often favor longer responses independent of quality, and conditioning on length removes this confounder from the $S\to Y$ mapping. Select the mode by OOF RMSE with a one-standard-error (1-SE) preference for monotone. Let $R^{\mathrm{OOF}}$ denote OOF predictions used along the IF path; the point estimate may use the pooled fit. The terminal isotonic step makes AutoCal-R mean-honest in either mode. We treat $\hat{f}$ as learned and propagate its uncertainty via OUA inference (Section 3.5).
Figure 3: AutoCal-R: monotone vs. two-stage calibration. Left: standard isotonic regression enforces monotonicity but cannot capture non-monotonic patterns in $\mathbb{E}[Y\mid S]$. Right: two-stage calibration (spline index $\to$ isotonic) can fit flexible patterns while preserving the mean. AutoCal-R automatically selects the mode via cross-validation.
3.2 Weight calibration (SIMCal-W: unit-mean ratios via OOF stacking of $S$-monotone candidates)

### v3

@@ 1 Introduction
Many teams now choose between system configurations (models, prompts, parameters, tools) by generating outputs and scoring them with an LLM judge. This is fast, and often wrong in a systematic way. Consider a common failure: you compared configuration A to B, the judge said A wins, you shipped A, and users hated it. The mechanism: verbosity and style bias. Judges frequently reward longer, more elaborate responses independent of actual quality; optimizing against the judge selected the wrong configuration. The standard workaround (reserve oracle labels for a small validation slice and hope the correlation holds) is ubiquitous in practice but statistically unsound.
CJE is a simple protocol: calibrate the judge to a small set of oracle labels, then audit whether that calibration remains valid for each target policy. When the audit fails, we do not “hope correlation holds”; we either recalibrate or refuse level claims.
Quick start for practitioners:
  • Use Direct + two-stage calibration. For open-ended generation under nontrivial policy shifts, logs-only OPE is impractical: target-typical trajectories receive little logger mass (low TTC), yielding a coverage-limited precision floor that weight stabilization cannot overcome.
  • Audit transport for high-stakes decisions (mean residual test per policy).
  • Gate on TTC $\geq 0.70$ before relying on logged data; in our benchmark, only near-clone policies pass this threshold.
@@
    Rankings can invert. Judge scores $S$ are treated as if they were rewards $Y$ on the same scale. They are not. Without calibration, higher $S$ can predict lower $Y$, inverting rankings entirely. This is the “surrogate paradox” in a new guise: surrogates can mislead catastrophically when the surrogate-target relationship shifts across contexts (Fleming & DeMets, 1996). The “You’re absolutely right!” phenomenon illustrates the cost: sycophantic responses score high on cheap proxies but low on actual user value. Teams know they should evaluate against long-term outcomes, but oracle labels are expensive, so they settle for proxies that may point the wrong direction.
@@
    Standard confidence intervals fail catastrophically. Naive CIs on uncalibrated judge scores achieve near-zero coverage: their 95% CIs capture the true value in 0/50 seeds. Bootstrap inference with bias-corrected estimation restores coverage to $\sim$}95% for Direct mode. (Figure 5: at 5% oracle fraction, calibration uncertainty contributes $\sim$}90% of total variance.)
@@
Why logs-only OPE fails for open-ended generation. Off-policy evaluation requires the logger to have visited regions where the target concentrates. In autoregressive text generation, the “action” is an entire response—a trajectory through a space of $|\mathcal{V}|^{T}$ possible token sequences. Two issues compound: (i) teacher-forcing propensities are empirically unreliable (even our clone policy achieves only 26% raw ESS vs. theoretical 100%), and (ii) for meaningfully different policies, TTC is 19–49%, indicating the logger rarely visits target-typical regions. Even with perfect propensities, CLE implies a hard precision floor when TTC is low. In constrained settings (short outputs, near-clone shifts, tool-choice bandits), OPE may remain viable; for open-ended generation under nontrivial policy shifts, Direct sidesteps these issues entirely.
@@
Recent work recognizes that LLM judges require calibration (Lee et al., 2025), and concurrent work (Chen et al., 2026) develops semiparametrically efficient, calibration-aware inference for mean parameters in the standard “large judge-labeled test + small human-labeled calibration” design. Guerdan et al. (2025) use doubly robust methods to address external validity when LLM judges are prompted with demographic personas. Our focus is complementary: we study multi-policy ranking under policy/deployment shift, introduce auditable transport diagnostics (policy-wise mean residual tests), and provide overlap/coverage diagnostics (CLE) and weight stabilization for off-policy evaluation, components not addressed by mean-parameter-only treatments. CJE fills this gap.
Auditable assumptions.
The core insight of CJE is to replace an uncheckable surrogacy assumption with a policy-wise moment restriction that we can audit. In causal inference, negative controls and proxy variables provide a template for this move: they turn untestable identification conditions into checkable diagnostics using auxiliary data (Lipsitch et al., 2010; Tchetgen Tchetgen et al., 2024). CJE applies this philosophy to LLM evaluation by budgeting a small oracle slice to test whether the judge-to-oracle calibration remains valid in each target context. The protocol is: calibrate on $\pi_{0}$, then audit transport on each $\pi^{\prime}$ (or time window, user segment, or task population) with a small oracle sample. Pass $\to$ reuse calibration; fail $\to$ recalibrate or refuse level claims for that cell. This transforms a brittle identification assumption into a standardized audit with an explicit decision rule.
@@
We introduce CJE, a framework that addresses all three failures. [^1] Library: pip install cje-eval (https://github.com/cimo-labs/cje); experiments: https://github.com/cimo-labs/cje-arena-experiments CJE casts offline evaluation as calibrated surrogate estimation with three core components:
  • Reward calibration: mean-preserving isotonic regression from judge score $S$ to oracle labels $Y$, with a two-stage option for covariate-dependent bias (e.g., response length).
  • Weight stabilization: variance-optimal stacking of monotone weight candidates, stabilizing importance weights for improved effective sample size.
  • Calibration-aware inference: bootstrap with bias correction that propagates calibration uncertainty into confidence intervals, achieving near-nominal coverage.
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019), learning $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, and applies it to LLM evaluation with auditable transport diagnostics. These components instantiate a single design rule: encode justified knowledge as model restrictions (closed convex sets) and project onto them. Building on established semiparametric theory, we show that such restrictions can only reduce variance (Theorem 2); with cross-fitting, our estimators attain the surrogate information bound. (Component names and technical details in Section 3.)
Beyond variance control, CJE includes a mean transport test that operationalizes the auditable assumption protocol: for each target policy $\pi^{\prime}$ or time period, we test whether the mean residual $\mathbb{E}[Y-f(S,X)]$ is zero. In our Arena benchmark, the base-trained calibration transports cleanly to clone, premium, and prompt-engineered policies, but fails for an adversarial “unhelpful” policy, where mean residuals reveal a 0.31-0.31 level shift. When the test fails, the failure is visible and actionable: either recalibrate with policy-specific data or refuse absolute value claims for that cell.
@@
  • OPE fails unexpectedly. We expected weight stabilization to enable OPE; it did not. Despite weight stabilization boosting ESS from $1\%$ to $80\%$, calibrated IPS remains near-random (47% pairwise). CLE explains why: high ESS is necessary but not sufficient when the logger rarely visits target-typical regions.
@@
    Methods: Mean-preserving reward calibration, weight stabilization for improved ESS, and calibration-aware bootstrap inference.
@@
    Theory: CJE adapts the surrogate index framework to LLM evaluation with a key innovation: transportability becomes auditable rather than assumed via the mean transport test. Inference is grounded in the efficient influence function for missing-outcome estimation, explaining why bootstrap with $\hat{\theta}_{\mathrm{aug}}$ achieves $\sim$}95% coverage where naive methods achieve 0%. The CLE bound provides a novel explanation for OPE failure despite high ESS, formalizing a precision floor under limited overlap.
@@
Readers primarily interested in deployment can focus on Section 3 (pipeline overview), Section 5.4 (seven practical rules), and Table 7 (estimator selection guide).
@@ 2 Background and Setup
Two operating modes: ranking vs. levels.
CJE supports two use cases with different requirements:
  • Ranking: Determine which policy is better ($V(\pi_{1})>V(\pi_{2})$?). Requires calibration to preserve order across policies. If calibration bias differs by policy (e.g., overestimates sycophantic responses), rankings can invert even without level claims.
  • Absolute levels: Report actual policy values ($V(\pi^{\prime})=0.72$) or make deployment decisions based on thresholds. Requires mean-unbiasedness: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$.
The transport audit catches both failure modes: policy-specific bias that inverts rankings and level shifts that bias absolute values. For low-stakes exploratory analysis, the audit may be deferred; for high-stakes decisions (deployment, cross-time comparisons, or when policy differences are small), run the audit.
@@
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019): learn a mapping $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, then use it for downstream estimation. For policy value estimation, we require only mean transport: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$, which we treat as auditable via a small oracle slice in each target context (4). We additionally impose monotonicity and mean-preservation as model restrictions. These are not identification assumptions, but deliberate regularization choices that stabilize calibration and prevent preference inversions.
@@
where $\sigma_{\mathcal{T}}^{2}:=\mathrm{ess\,inf}_{(x,a)\in\mathcal{T}}\mathrm{Var}(Y\mid X{=}x,A{=}a)$ is the minimal outcome noise in $\mathcal{T}$, and $\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})$ measures shape mismatch inside $\mathcal{T}$ (proof in Appendix C.12). The bound has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions; (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, inflating the floor even with good coverage; and (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
@@
Automatic judges (LLM-as-judge or preference models) provide scalable scoring (Ouyang et al., 2022; Bai et al., 2022; Zheng et al., 2023; Kim et al., 2024; Kocmi & Federmann, 2023) but are correlational and may drift (Dietz et al., 2025). Viewing $S$ as a surrogate connects to surrogate endpoints and mediation (Prentice, 1989; Robins & Greenland, 1992; Frangakis & Rubin, 2002; Pearl, 2012; VanderWeele, 2015). The surrogate paradox, wherein a policy improves $S$ and $S$ correlates with $Y$ yet the policy harms $Y$, can arise even with high correlation (VanderWeele, 2013); our transport test (4) provides a necessary check. Calibrating $R=f(S)$ enables unbiased estimation of $V(\pi^{\prime})$ when the transport condition holds: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ (4). Mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=\mu(S)$) is one sufficient condition guaranteeing transport; in CJE, level claims are gated by a per-policy transport audit. Without that audit, transport must be assumed. The surrogate also supplies a one-dimensional index that stabilizes weights.
@@
Figure 1: Judge-as-surrogate under policy and environment shift. $X$: prompt; AA: response; $S$: judge score; $Y$: oracle label; $\pi$: policy; $E$: environment. $S$ and $Y$ are parallel measurements of the response. Solid arrows: causal structure ($\pi\to A$: policy determines response). Dashed red arrows: mechanism-shift edges indicating the $S$-$Y$ calibration may differ across policies or environments; these are not causal effects of $\pi$ or $E$ on $Y$ holding $(X,A)$ fixed. Plain English: Different policies generate different response styles, and the judge may be biased toward certain styles (e.g., verbosity). CJE’s transport test catches this via $\mathbb{E}[Y-f(S,X)]=0$.
@@
Transport Failure $\neq$ Distribution Shift
Distribution shift: The prompts change between training and deployment.
Transport failure: Prompts can be identical, but your policy changes the kinds of responses it generates, and the judge’s bias interacts with that (style, verbosity, self-preference). The mapping $S\to Y$ changes even though $X$ doesn’t. CJE tests this directly: for each target policy $\pi^{\prime}$, we check whether $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$. If not, the calibration doesn’t transport—either recalibrate or refuse level claims.
Calibration-aware uncertainty & influence-function stacking.
Treating learned $R=\hat{f}(S)$ as fixed understates uncertainty when the oracle slice is small. CJE addresses this via calibration-aware inference, which propagates calibration uncertainty into confidence intervals using a delete-one-fold jackknife (Efron & Stein, 1981; Bickel et al., 1993). For ensemble estimation, CJE stacks multiple estimators by minimizing IF covariance over the simplex (Wolpert, 1992; van der Laan et al., 2007). Details in Section 3.
@@ 3 Methods
The CJE pipeline has three components: reward calibration to prevent preference inversion, weight stabilization for importance-weighted estimators, and uncertainty-aware inference for valid CI coverage. For open-ended generation under nontrivial policy shifts, reward calibration and inference are the essential components; weight stabilization addresses IPS/DR variance but faces coverage limitations when TTC is low (Section 1). Figure 2 provides an overview.
@@
CJE encodes justified knowledge as model restrictions: (i) monotonicity and mean-preservation for reward calibration, (ii) mean-one constraint with $S$-monotonicity for weights, (iii) nuisance-orthogonality for estimators, and (iv) convex combination for stacking. All learners are cross-fitted; by the Efficiency via Model Restriction theorem, these restrictions lower the efficiency bound (see Section 4).
3.1 Reward calibration (two-stage isotonic)
@@
  • Two-stage mode (recommended): Fit a smooth index $Z(S,X)=g(S,X)$ (splines+ridge), map to mid-ranks $U=\mathrm{ECDF}\{Z(S,X)\}$, then fit isotonic $\hat{h}_{\uparrow}(U)$. Predictions are $R=\hat{h}_{\uparrow}(\mathrm{ECDF}\{g(S,X)\})$.
  • Monotone-only mode: Isotonic regression directly on $S$: $\hat{f}_{\uparrow}\in\arg\min_{f\in\mathcal{M}_{\uparrow}}\sum_{i\in O}\big(Y_{i}-f(S_{i})\big)^{2}$. PAVA preserves the slice mean exactly. Use when covariates $X$ are unavailable.
Remark (covariate flexibility). The terminal isotonic step preserves mean-honesty regardless of the first-stage model, so $g$ can be any cross-fitted learner (splines, random forests, gradient boosting) and can incorporate covariates $X$ beyond $S$. In our experiments, including response length as a first-stage covariate improves ranking accuracy: LLM judges often favor longer responses independent of quality, and conditioning on length removes this confounder from the $S\to Y$ mapping. Let $R^{\mathrm{OOF}}$ denote OOF predictions used along the IF path; the point estimate may use the pooled fit. The terminal isotonic step makes the calibrator mean-honest in either mode. We treat $\hat{f}$ as learned and propagate its uncertainty via calibration-aware inference (Section 3.5).
Figure 3: Reward calibration: monotone vs. two-stage. Left: monotone-only isotonic regression enforces monotonicity but cannot capture non-monotonic patterns in $\mathbb{E}[Y\mid S]$. Right: two-stage calibration (spline index $\to$ isotonic) can fit flexible patterns while preserving the mean.
3.2 Weight stabilization (unit-mean ratios via OOF stacking of $S$-monotone candidates)
Scope and practical relevance: Weight stabilization applies only to IPS and DR estimators. For open-ended generation under nontrivial policy shifts, coverage sparsity (low TTC) limits the utility of logged data (Section 1); in such regimes, Direct is preferred. Weight stabilization remains valuable in constrained settings (short outputs, near-clone shifts, tool-choice bandits) where TTC $\geq 0.70$, or when combining logged data with fresh draws via DR. The Direct method—which generates fresh responses under $\pi^{\prime}$—requires no importance weights.